Tips for a Succesful Presentation ILAS 2008, Cancún

The Local Organizing Committee would like to respectfully offer the following points for a successful presentation of your talk at ILAS 2008.

• There will be a laptop with a projector for your presentation. Make sure you bring your presentation on a CD or a USB flash memory. The laptops have no floppy drive and no internet connection.

• The laptops will contain Microsoft Power Point and Adobe Acrobat Reader. Make sure your presentation runs with one of these softwares. A PDF presentation generated using L^{A}TEX with the Beamer class seems to be the popular choice these days! Other useful L^{A}TEXpackages are foiltex and seminar.

• There will be no overhead projectors nor whiteboards/blackboards available.

• Please make sure that the font size is adequate for a large room. Any size larger than 24pt should be adequate. Also, make sure that the pages appear on a landscape orientation.

• Audiences do not appreciate computer presentations that consist of pages of a paper, published or to be published. Please avoid doing this.

• The pages of your presentation should not be overcrowded with lines. About 10 lines per page is usually adequate. Try not to put more than 15 lines!!

• Try to keep your presentation short. A good rule of thumb is: a bit less than one page per minute of presentation, although this varies with each case. Practice your presentation to see how long it takes.

• If you must include proofs, try to make them as short and intelligible as you can.

• Avoid long calculations. All calculations included should be kept short enough to fit in the time alloted for that page.

• Remember that you have been an audience member. Try to avoid doing those things that have bothered you when you have listened to a talk.

• Above all, make your experience giving a talk a pleasant one. Have fun in Cancún!